# Homotopy connectivity of Čech complexes of spheres

### v1

@@ 1. Introduction
What are the homotopy types of intrinsic Čech complexes of spheres? Equip the $n$-sphere $S^{n}$ with the geodesic metric of diameter $\pi$. If the scale $r$ satisfies $0<r\leq\frac{\pi}{2}$, then the open balls $\{B_{S^{n}}(x;r)\}_{x\in S^{n}}$ form a good cover of the sphere $S^{n}$, meaning that any intersection of balls is either empty or contractible. Since $S^{n}$ is paracompact as it is a metric space, the nerve lemma implies that the intrinsic Čech complex $\mathrm{\check{C}}(S^{n};r)\coloneqq\mathrm{Nerve}(\{B_{S^{n}}(x;r)\}_{x\in S^{n}})$ is homotopy equivalent to $S^{n}$ ([22, Theorem 4G.3]; see also [43, Theorem 13.1.3]). But what are the homotopy types of $\mathrm{\check{C}}(S^{n};r)$ when $\frac{\pi}{2}<r<\pi$?
The case $n=1$ is well-understood by [1, Section 9]: as the scale $r$ increases, the complexes $\mathrm{\check{C}}(S^{1};r)$ obtain the homotopy types of $S^{1}$, $S^{3}$, $S^{5}$, $S^{7}$, …. We also refer the reader to [2] for a description of the homotopy type of $\mathrm{\check{C}}(X;r)$ for $X$ an arbitrary finite subset of $S^{1}$, and to [28] for a description of the expected Euler characteristic of $\mathrm{\check{C}}(X;r)$ for $X$ a finite subset chosen uniformly at random from $S^{1}$. However, for $n\geq 2$, essentially nothing is known about the homotopy types of $\mathrm{\check{C}}(S^{n};r)$ for $\frac{\pi}{2}<r<\pi$, except that they are simply connected by [45, Theorem 11.2]. In particular, though the first new homotopy type beyond $S^{n}$ (i.e., the homotopy type of $\mathrm{\check{C}}(S^{n};\frac{\pi}{2}+\varepsilon)$ for $\varepsilon>0$ sufficiently small) seems within reach, to our knowledge no conjecture for this homotopy type has appeared in the literature. Can we bound the homotopy connectivity of $\mathrm{\check{C}}(S^{n};r)$ as a function of the scale $r$?
@@
Figure 1. Intervals where $\mathrm{\check{C}}(S^{2};r)$ may have connectivity $k-1$, given by Theorem 1.1. The blue endpoints are plotted using only approximate values of $\operatorname{cov}_{S^{2}}(2k+2)$ or of $2\operatorname{cov}_{S^{2}}(k+1)$; see [41].
@@
Let $r=\pi-\delta$ for $\delta\in(0,\pi)$. Note that $\operatorname{numCover}_{S^{n}}(\delta)\to\infty$ as $\delta\to 0$. So, from the left-hand inequality in Corollary 1.2, we see that as $\delta$ tends to 00, the connectivity of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ tends to infinity. On the other hand, since $\operatorname{numCover}_{S^{n}}(\delta)<\infty$ for all $\delta>0$, we get from the right-hand inequality in Corollary 1.2 that the connectivity of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is always finite. Hence, as $r$ varies between 00 and $\pi$, $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))$ attains infinitely many integer values. ∎
@@
How good are our bounds on $\mathrm{\check{C}}(S^{n};\pi-\delta)$? We are able to test this in the case of $n=1$. From [1, Theorem 9.8], we know that $\mathrm{\check{C}}(S^{1};r)$ obtains the homotopy types $S^{1},S^{3},S^{5},S^{7},\ldots$ as $r$ increases. Thus, we compare the scales obtained in Theorem 1.1 with the actual scales of those connectivities, and it turns out that for $n=1$, the lower bound on $\delta$ in the theorem is off by approximately a multiplicative factor of $4$, whereas the upper bound on $\delta$ is tight. See Figure 2.
Figure 2. The homotopy types of $\mathrm{\check{C}}(S^{1};r)$ as $r$ varies [1] are indicated by black bars. Theorem 1.1 gives intervals where $\mathrm{\check{C}}(S^{1};r)$ may have connectivity $k-1$, which are indicated by colored bars. Left endpoints of orange bars (when $k-1$ is even) are tight.
@@ 2. Related work
This recovers the classical result that the chromatic number of the Borsuk graph is at least $n+2$ for scales $\delta$ which are only slightly less than $\pi$ [29, 37, 38] [^1] In these references, the result for chromatic numbers is proven for closed Borsuk graphs $\mathrm{Bor}[S^{n};\delta]$, which have an edge between vertices $v$ and $w$ whenever $d(v,w)\geq\delta$. However, this result remains true (see Remark 3.1) when Borsuk graphs are considered with the “open” convention [19] used in this paper., since in this regime, the nerve lemma [13, 22] applies to give $\mathrm{\check{C}}(S^{n};\pi-\delta)\simeq S^{n}$, yielding $\chi(\mathrm{Bor}(S^{n};\delta))\geq\operatorname{conn}(S^{n})+3=n+2$. There is furthermore an equality $\chi(\mathrm{Bor}(S^{n};\delta))=n+2$ for $\delta$ sufficiently close to $\pi$. However, in this paper, we allow the scale $\delta$ to be far away from $\pi$. Using known bounds on the chromatic numbers of Borusk graphs in terms of packings and coverings of spheres (see, for example, [36]), we obtain upper bounds on the connectivity of Čech complexes of spheres at all scale parameters.
@@
Researchers in applied topology use Čech and Vietoris–Rips complexes to approximate the shape of a dataset. If one’s dataset is sampled from a manifold, then in the limit, as more and more points are sampled, the persistent homology of the sample converges to the persistent homology of the manifold. For small scales, Hausmann’s theorem [23] implies that the Vietoris–Rips complex of the manifold is homotopy equivalent to the manifold, and the nerve lemma [13] implies that the intrinsic Čech complex of the manifold is homotopy equivalent to the manifold. However, we do not yet have a good mathematical understanding of the behavior of the Vietoris–Rips or Čech complexes of manifolds at larger scales, even though the whole idea of persistent homology is to allow the scale to vary from small to large. The homotopy types of Vietoris–Rips complexes of spheres is a difficult but largely open problem [1, 26, 5], which has important connections to Gromov–Hausdorff distances between unit spheres of different dimensions [27, 4, 21]. Though the homotopy types of intrinsic Čech complexes of spheres have been studied less, we think this may be an even more natural problem. Indeed, these homotopy types can be viewed as generalizations of the nerve lemma, for one may ask: What are the homotopy types of the nerve of balls in a sphere (or more generally in a manifold) when the radii are now large enough so that the hypotheses of the nerve lemma are no longer satisfied?
We remark that the homotopy types of the intrinsic Čech complexes of the circle are fully understood by [1, Theorem 9.8]. Indeed, we have $\mathrm{\check{C}}(S^{1};r)\simeq S^{2k+1}$ for $\frac{\pi k}{k+1}<r\leq\frac{\pi(k+1)}{k+2}$ for $k\in{\mathbb{N}}$. The way to interpret these homotopy types is as follows. For $0<r\leq\frac{\pi}{2}$, open balls of radius $r$ (circular arcs of length $2r$) in the circle are geodesically convex, and hence, the nerve lemma applies to give $\mathrm{\check{C}}(S^{1};r)\simeq S^{1}$. However, once $r$ exceeds $\frac{\pi}{2}$, a ball of radius $r$ is no longer geodesically convex, and indeed, the intersection of two antipodal balls is disconnected. Therefore, the hypotheses of the nerve lemma are not satisfied. So, up to homotopy type, a circle’s worth of disks are attached, giving $\mathrm{\check{C}}(S^{1};r)\simeq S^{1}*S^{1}=S^{3}$ for $\frac{\pi}{2}<r\leq\frac{2\pi}{3}$. Once $r$ exceeds $\frac{2\pi}{3}$, the intersection of three evenly spaced balls now has three connected components, and the homotopy type becomes $(S^{1}*S^{1})*S^{1}=S^{3}*S^{1}=S^{5}$. More generally, for $\frac{\pi k}{k+1}<r\leq\frac{\pi(k+1)}{k+2}$, we have that $\mathrm{\check{C}}(S^{1};r)$ is homotopy equivalent to a $(k+1)$-fold join of circles $S^{1}*\cdots*S^{1}=S^{2k+1}$.
@@ 3.1. Metric spaces
Let $(X,d)$ be a metric space. Further, let $B(x,r)$ be the open ball of radius $r$ centered at $x$, i.e., $B(x;r)=\{y\in X\mid d(y,x)<r\}$. Also, let $B[x,r]=\{y\in X\mid d(y,x)\leq r\}$ be the closed ball of radius $r$ centered at $x$.
@@
[…]
@@
$\operatorname{numPack}_{X}(\delta)=\max\left\{m\geq 1\hskip 2.84526pt\middle|\hskip 2.84526pt\text{ there exist }x_{1},\ldots,x_{m}\in X\text{ such that }d(x_{i},x_{j})>\delta\text{ for all }i,j\right\}.$
@@ 3.3. Simplicial complexes
A simplicial complex $K$ on a vertex set $V$ is a collection of subsets of $V$, containing all singletons, such that if $\sigma\in K$ and $\tau\subseteq\sigma$, then $\tau\in K$ as well. In this paper, we identify simplicial complexes with their geometric realizations. We will be most interested in simplicial complexes that can be built on top of metric spaces, such as Čech and Vietoris–Rips simplicial complexes.
@@
$\mathrm{\check{C}}(X;r)=\left\{\text{finite }\sigma\subseteq X~{}|~{}\bigcap_{x\in\sigma}B(x,r)\neq\emptyset\right\}.$
@@ 3.4. Graphs
While Borsuk graphs were originally introduced in [19] with the open ($>$) convention that we use in this paper, Borsuk graphs with the closed ($\geq$) convention have been more popular in the literature since; see, for example, [30, 29]. The closed Borsuk graph of $S^{n}$ for scale $\delta>0$, denoted $\mathrm{Bor}[S^{n};\delta]$, is a simple graph whose vertex set is $S^{n}$ and whose edges are all pairs of vertices $x,y\in S^{n}$ with $d(x,y)\geq\delta$. For any small $\varepsilon>0$, we have that
@@ 4. A lower bound on connectivity
Let $\varepsilon>0$ be such that $\delta+\varepsilon<\operatorname{cov}_{S^{n}}(2k+2)$. Thus, no $22k+$ 2 closed balls of radius $\delta+\varepsilon$ can cover $S^{n}$. Let $X$ be a finite subset of $S^{n}$ that is $\varepsilon$-dense, i.e., such that any open $\varepsilon$-ball in $S^{n}$ contains a point of $X$. Then for any collection of $22k+$ 2 points $x_{1},\ldots,x_{2k+2}\in X$, we have by the definition of $\operatorname{cov}_{S^{n}}(2k+2)$ that
@@
Since $X$ is $\varepsilon$-dense, there is some point $x\in X$ with $x\in\cap_{i=1}^{2k+2}B_{X}(x_{i};\pi-\delta)$. Thus, $\mathrm{\check{C}}(X;\pi-\delta)$ is $(2k+2)$-conic. Since $X$ is finite, we use Barmak’s theorem [20] to conclude that $\operatorname{conn}(\mathrm{\check{C}}(X;\pi-\delta))\geq k$.
Now, for any $i\leq k$, let us consider a continuous map $f\colon S^{i}\to\mathrm{\check{C}}(S^{n};\pi-\delta)$. To show that $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k$, we show that $f$ is null-homotopic, and hence $\pi_{i}(\mathrm{\check{C}}(S^{n};\pi-\delta))$ is trivial for $i\leq k$. We first note that the image $\text{Im}(f)$ is compact since $S^{i}$ is compact. Thus, there exists a subcomplex $\mathrm{\check{C}}(X;\pi-\delta)\subseteq\mathrm{\check{C}}(S^{n};\pi-\delta)$ with a finite vertex set $X$ such that $\text{Im}(f)\subseteq\mathrm{\check{C}}(X;\pi-\delta)$. Without loss of generality, $X$ can be assumed to be $\varepsilon$-dense in $S^{n}$ because, if needed, finitely many more vertices can be added to make $X$ sufficiently dense. Since $\operatorname{conn}(\mathrm{\check{C}}(X;\pi-\delta))\geq k$, we have $\pi_{i}(\mathrm{\check{C}}(X;\pi-\delta))=0$ for all $i\leq k$, and thus $f$ is null-homotopic in $\mathrm{\check{C}}(X;\pi-\delta)$. Let $H\colon S^{i}\times I\to\mathrm{\check{C}}(X;\pi-\delta)$ be a null-homotopy for $f$. If $j\colon\mathrm{\check{C}}(X;\pi-\delta)\hookrightarrow\mathrm{\check{C}}(S^{n};\pi-\delta)$ is the inclusion induced by the inclusion $X\hookrightarrow S^{n}$, then the composition
@@
Our proof technique showing that $\mathrm{\check{C}}(X;\pi-\delta)$ is $(2k+2)$-conic in fact shows something stronger, namely that any subcomplex $L$ of $\mathrm{\check{C}}(X;\pi-\delta)$ with at most $22k+$ 2 vertices is contained in a simplex. Does this mean that there is potentially room for improvement to prove a stronger result than Proposition 4.1?

### v2

@@ 1. Introduction
What are the homotopy types of intrinsic Čech complexes of spheres? Equip the $n$-sphere $S^{n}$ with the geodesic metric of diameter $\pi$. If the scale $r$ satisfies $0<r\leq\frac{\pi}{2}$, then the open balls $\{B_{S^{n}}(x;r)\}_{x\in S^{n}}$ form a good cover of the sphere $S^{n}$, meaning that any intersection of balls is either empty or contractible. Since $S^{n}$ is paracompact as it is a metric space, the nerve lemma implies that the intrinsic Čech complex $\mathrm{\check{C}}(S^{n};r)\coloneqq\mathrm{Nerve}(\{B_{S^{n}}(x;r)\}_{x\in S^{n}})$ is homotopy equivalent to $S^{n}$ ([22, Theorem 4G.3]; see also [44, Theorem 13.1.3]). But what are the homotopy types of $\mathrm{\check{C}}(S^{n};r)$ when $\frac{\pi}{2}<r<\pi$?
The case $n=1$ is well-understood by [2, Section 9]: as the scale $r$ increases, the complexes $\mathrm{\check{C}}(S^{1};r)$ obtain the homotopy types of $S^{1}$, $S^{3}$, $S^{5}$, $S^{7}$, …. We also refer the reader to [1] for a description of the homotopy type of $\mathrm{\check{C}}(X;r)$ for $X$ an arbitrary finite subset of $S^{1}$, and to [28] for a description of the expected Euler characteristic of $\mathrm{\check{C}}(X;r)$ for $X$ a finite subset chosen uniformly at random from $S^{1}$. However, for $n\geq 2$, essentially nothing is known about the homotopy types of $\mathrm{\check{C}}(S^{n};r)$ for $\frac{\pi}{2}<r<\pi$, except that they are simply connected by [46, Theorem 11.2]. In particular, though the first new homotopy type beyond $S^{n}$ (i.e., the homotopy type of $\mathrm{\check{C}}(S^{n};\frac{\pi}{2}+\varepsilon)$ for $\varepsilon>0$ sufficiently small) seems within reach, to our knowledge no conjecture for this homotopy type has appeared in the literature. Can we bound the homotopy connectivity of $\mathrm{\check{C}}(S^{n};r)$ as a function of the scale $r$?
@@
Figure 1. Intervals where $\mathrm{\check{C}}(S^{2};r)$ may have connectivity k1k-1, given by Theorem 1.1. The blue endpoints are plotted using only approximate values of $\operatorname{cov}_{S^{2}}(2k+2)$ or of $2\operatorname{cov}_{S^{2}}(k+1)$; see [42].
@@
Let $r=\pi-\delta$ for $\delta\in(0,\pi)$. Note that $\operatorname{numCover}_{S^{n}}(\delta)\to\infty$ as $\delta\to 0$. So, from the left-hand inequality in Corollary 1.2, we see that as $\delta$ tends to 0, the connectivity of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ tends to infinity. On the other hand, since $\operatorname{numCover}_{S^{n}}(\delta)<\infty$ for all $\delta>0$, we get from the right-hand inequality in Corollary 1.2 that the connectivity of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is always finite. Hence, as $r$ varies between 0 and $\pi$, $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))$ attains infinitely many integer values. ∎
@@
How good are our bounds on $\mathrm{\check{C}}(S^{n};\pi-\delta)$? We are able to test this in the case of $n=1$. From [2, Theorem 9.8], we know that $\mathrm{\check{C}}(S^{1};r)$ obtains the homotopy types $S^{1},S^{3},S^{5},S^{7},\ldots$ as $r$ increases. Thus, we compare the scales obtained in Theorem 1.1 with the actual scales of those connectivities, and it turns out that for $n=1$, the lower bound on $\delta$ in the theorem is off by approximately a multiplicative factor of $4$, whereas the upper bound on $\delta$ is tight. See Figure 2.
Figure 2. The homotopy types of $\mathrm{\check{C}}(S^{1};r)$ as $r$ varies [2] are indicated by black bars. Theorem 1.1 gives intervals where $\mathrm{\check{C}}(S^{1};r)$ may have connectivity k1k-1, which are indicated by colored bars. Left endpoints of orange bars (when k1k-1 is even) are tight.
@@ 2. Related work
This recovers the classical result that the chromatic number of the Borsuk graph is at least $n+2$ for scales $\delta$ which are only slightly less than $\pi$ [29, 38, 39] [^1] In these references, the result for chromatic numbers is proven for closed Borsuk graphs $\mathrm{Bor}[S^{n};\delta]$, which have an edge between vertices $v$ and $w$ whenever $d(v,w)\geq\delta$. However, this result remains true (see Remark 3.1) when Borsuk graphs are considered with the “open” convention [19] used in this paper., since in this regime, the nerve lemma [13, 22] applies to give $\mathrm{\check{C}}(S^{n};\pi-\delta)\simeq S^{n}$, yielding $\chi(\mathrm{Bor}(S^{n};\delta))\geq\operatorname{conn}(S^{n})+3=n+2$. There is furthermore an equality $\chi(\mathrm{Bor}(S^{n};\delta))=n+2$ for $\delta$ sufficiently close to $\pi$. However, in this paper, we allow the scale $\delta$ to be far away from $\pi$. Using known bounds on the chromatic numbers of Borsuk graphs in terms of packings and coverings of spheres (see, for example, [37]), we obtain upper bounds on the connectivity of Čech complexes of spheres at all scale parameters.
@@
Researchers in applied topology use Čech and Vietoris–Rips complexes to approximate the shape of a dataset. If one’s dataset is sampled from a manifold, then in the limit, as more and more points are sampled, the persistent homology of the sample converges to the persistent homology of the manifold. For small scales, Hausmann’s theorem [23] implies that the Vietoris–Rips complex of the manifold is homotopy equivalent to the manifold, and the nerve lemma [13] implies that the intrinsic Čech complex of the manifold is homotopy equivalent to the manifold. However, we do not yet have a good mathematical understanding of the behavior of the Vietoris–Rips or Čech complexes of manifolds at larger scales, even though the whole idea of persistent homology is to allow the scale to vary from small to large. Determining the homotopy types of Vietoris–Rips complexes of spheres is a difficult but largely open problem [2, 26, 5], which has important connections to Gromov–Hausdorff distances between unit spheres of different dimensions [27, 4, 21]. Though the homotopy types of intrinsic Čech complexes of spheres have been studied less, we think this may be an even more natural problem. Indeed, these homotopy types can be viewed as generalizations of the nerve lemma, for one may ask: What are the homotopy types of the nerve of balls in a sphere (or more generally in a manifold) when the radii are now large enough so that the hypotheses of the nerve lemma are no longer satisfied?
We remark that the homotopy types of the intrinsic Čech complexes of the circle are fully understood by [2, Theorem 9.8]. Indeed, we have $\mathrm{\check{C}}(S^{1};r)\simeq S^{2k+1}$ for $\frac{\pi k}{k+1}<r\leq\frac{\pi(k+1)}{k+2}$ for $k\in{\mathbb{N}}$. The way to interpret these homotopy types is as follows. For $0<r\leq\frac{\pi}{2}$, open balls of radius $r$ (circular arcs of length $2r$) in the circle are geodesically convex, and hence, the nerve lemma applies to give $\mathrm{\check{C}}(S^{1};r)\simeq S^{1}$. However, once $r$ exceeds $\frac{\pi}{2}$, a ball of radius $r$ is no longer geodesically convex, and indeed, the intersection of two antipodal balls is disconnected. Therefore, the hypotheses of the nerve lemma are not satisfied. So, up to homotopy type, a circle’s worth of disks are attached, giving $\mathrm{\check{C}}(S^{1};r)\simeq S^{1}*S^{1}=S^{3}$ for $\frac{\pi}{2}<r\leq\frac{2\pi}{3}$. Once $r$ exceeds $\frac{2\pi}{3}$, the intersection of three evenly spaced balls now has three connected components, and the homotopy type becomes $(S^{1}*S^{1})*S^{1}=S^{3}*S^{1}=S^{5}$. More generally, for $\frac{\pi k}{k+1}<r\leq\frac{\pi(k+1)}{k+2}$, we have that $\mathrm{\check{C}}(S^{1};r)$ is homotopy equivalent to a $(k+1)$-fold join of circles $S^{1}*\cdots*S^{1}=S^{2k+1}$.
@@ 3.1. Metric spaces
Let $(X,d)$ be a metric space. Further, let $B(x,r)$ be the open ball of radius $r$ centered at $x$, i.e., $B(x;r)=\{y\in X\mid d(y,x)<r\}$. Also, let $B[x,r]=\{y\in X\mid d(y,x)\leq r\}$ be the closed ball of radius $r$ centered at $x$. On occasion, we may write $B_{X}(x,r)$ or $B_{X}[x,r]$, for clarity.
@@
$\operatorname{cov}_{X}(k)=\inf\left\{r\geq 0\hskip 2.84526pt\middle|\hskip 2.84526pt\text{ there exist }x_{1},\ldots,x_{k}\in X\text{ such that }\bigcup_{i=1}^{k}B[x_{i},r]=X\right\}.$
@@
[…]
@@ 3.3. Simplicial complexes
A simplicial complex $K$ on a vertex set $V$ is a collection of finite subsets of $V$, containing all singletons, such that if $\sigma\in K$ and $\tau\subseteq\sigma$, then $\tau\in K$ as well. In this paper, we identify simplicial complexes with their geometric realizations. We will be most interested in simplicial complexes that can be built on top of metric spaces, such as Čech and Vietoris–Rips simplicial complexes.
@@
$\mathrm{\check{C}}(X;r)=\left\{\text{finite }\sigma\subseteq X\ \middle|\ \bigcap_{x\in\sigma}B(x,r)\neq\emptyset\right\}.$
@@ 3.4. Graphs
While Borsuk graphs were originally introduced in [19] with the open (>>) convention that we use in this paper, Borsuk graphs with the closed ($\geq$) convention have been more popular in the literature since; see, for example, [30, 29]. The closed Borsuk graph of $S^{n}$ for scale $\delta>0$, denoted $\mathrm{Bor}[S^{n};\delta]$, is a simple graph whose vertex set is $S^{n}$ and whose edges are all pairs of vertices $x,y\in S^{n}$ with $d(x,y)\geq\delta$. For any small $\varepsilon>0$, we have that
@@ 4. A lower bound on connectivity
Let $\varepsilon>0$ be such that $\delta+\varepsilon<\operatorname{cov}_{S^{n}}(2k+2)$. Thus, no $2k+2$ closed balls of radius $\delta+\varepsilon$ can cover $S^{n}$. Let $X$ be a finite subset of $S^{n}$ that is $\varepsilon$-dense, i.e., such that any open $\varepsilon$-ball in $S^{n}$ contains a point of $X$. Then for any collection of $2k+2$ points $x_{1},\ldots,x_{2k+2}\in X$, we have by the definition of $\operatorname{cov}_{S^{n}}(2k+2)$ that
@@
Since $X$ is $\varepsilon$-dense, there is some point $x\in X$ with $x\in\cap_{i=1}^{2k+2}B_{X}(x_{i};\pi-\delta)$, where we now consider balls in $X$ instead of balls in $S^{n}$. Thus, $\mathrm{\check{C}}(X;\pi-\delta)$ is $(2k+2)$-conic. Since $X$ is finite, we use Barmak’s theorem [20] to conclude that $\operatorname{conn}(\mathrm{\check{C}}(X;\pi-\delta))\geq k$.
Now, for any $i\leq k$, let us consider a continuous map $f\colon S^{i}\to\mathrm{\check{C}}(S^{n};\pi-\delta)$. To show that $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k$, we show that $f$ is null-homotopic, and hence $\pi_{i}(\mathrm{\check{C}}(S^{n};\pi-\delta))$ is trivial for $i\leq k$. We first note that the image $\text{Im}(f)$ is compact since $S^{i}$ is compact. Thus, there exists a subcomplex $\mathrm{\check{C}}(X;\pi-\delta)\subseteq\mathrm{\check{C}}(S^{n};\pi-\delta)$ with a finite vertex set $X$ such that $\text{Im}(f)\subseteq\mathrm{\check{C}}(X;\pi-\delta)$; see, for instance, [22, Proposition A.1]. Without loss of generality, $X$ can be assumed to be $\varepsilon$-dense in $S^{n}$ because, if needed, finitely many more vertices can be added to make $X$ sufficiently dense. Since $\operatorname{conn}(\mathrm{\check{C}}(X;\pi-\delta))\geq k$, we have $\pi_{i}(\mathrm{\check{C}}(X;\pi-\delta))=0$ for all $i\leq k$, and thus $f$ is null-homotopic in $\mathrm{\check{C}}(X;\pi-\delta)$. Let $H\colon S^{i}\times I\to\mathrm{\check{C}}(X;\pi-\delta)$ be a null-homotopy for $f$. If $j\colon\mathrm{\check{C}}(X;\pi-\delta)\hookrightarrow\mathrm{\check{C}}(S^{n};\pi-\delta)$ is the inclusion induced by the inclusion $X\hookrightarrow S^{n}$, then the composition
@@
Our proof technique showing that $\mathrm{\check{C}}(X;\pi-\delta)$ is $(2k+2)$-conic in fact shows something stronger, namely that any subcomplex $L$ of $\mathrm{\check{C}}(X;\pi-\delta)$ with at most $2k+2$ vertices is contained in a simplex. Does this mean that there is potentially room for improvement to prove a stronger result than Proposition 4.1?